# Cost-aware Stopping for Bayesian Optimization

### v1

@@ C.3 Additional Experiment Results: Empirical
As discussed in Section 3.2, there are two variants of the LogEIPC acquisition functions and corresponding stopping rules under the unknown-cost setting. We refer to the variant using (13) as LogEIPC-inv, and the one using (14) as LogEIPC-exp. Recall that LogEIPC-exp is equivalent to PBGI as a stopping rule (though not equivalent as an acquisition function), while LogEIPC-inv differs from both not only as an acquisition function but also as a stopping rule. In Figure 17, we show the performance of acquisition function and stopping rule combinations under the unknown-cost setting, using LogEIPC-inv for the LogEIPC acquisition function. The results are qualitatively similar to the known-cost setting.
Figure 17: Comparison of cost-adjusted simple regret across acquisition function and stopping rule combinations under the unknown-cost setting on LCBench. Results are qualitatively similar to those in the known-cost setting.
C.4 Additional Experiment Results: Bayesian Regret
@@
In this section, we present the complete Bayesian regret results. Figures 18, 19 and 20 show the 1D experiments, and Figures 21, 22 and 23 show the 8D experiments. Each figure corresponds to one cost setting (uniform, linear or periodic) and three values of the cost-scaling factor, $\lambda=10^{-1},10^{-2},10^{-3}$. In all of the experimental results, we observe that PBGI/LogEIPC acquisition function + PBGI/LogEIPC stopping achieves cost-adjusted regret that is not only competitive with the baselines, but is also competitive regarding the best in hindsight fixed iteration stopping and often competitive even comparing to hindsight optimal stopping. These results indicate that our automatic stopping rule can replace manual selection of stopping times without loss in performance. Figures 18, 19, 20, 21, 22 and 23 also show that our PBGI/LogEIPC stopping rule outperforms other baselines when the cost-scaling factor $\lambda$ is large, indicating that it’s an especially suitable stopping criteria when evaluation is expensive.
@@
Figure 18: Comparison of cost-adjusted simple regret across acquisition function and stopping rule combinations in the 1D Bayesian regret experiments, with cost-scaling factor $\lambda=10^{-1},10^{-2},10^{-3}$ and under uniform cost. The dashed line in each subplot represent fixed iteration stopping (e.g., the y-axis value of the line at iteration $50$ represent the cost-adjusted regret when always stopping at iteration $50$).
@@
Figure 19: Comparison of cost-adjusted simple regret across acquisition function and stopping rule combinations in the 1D Bayesian regret experiments, with cost-scaling factor $\lambda=10^{-1},10^{-2},10^{-3}$ and under linear cost. The dashed line in each subplot represent fixed iteration stopping (e.g., the y-axis value of the line at iteration $50$ represent the cost-adjusted regret when always stopping at iteration $50$).
@@
Figure 20: Comparison of cost-adjusted simple regret across acquisition function and stopping rule combinations in the 1D Bayesian regret experiments, with cost-scaling factor $\lambda=10^{-1},10^{-2},10^{-3}$ and under periodic cost. The dashed line in each subplot represent fixed iteration stopping (e.g., the y-axis value of the line at iteration $50$ represent the cost-adjusted regret when always stopping at iteration $50$).
@@
Figure 21: Comparison of cost-adjusted simple regret across acquisition function and stopping rule combinations in the 8D Bayesian regret experiments, with cost-scaling factor $\lambda=10^{-1},10^{-2},10^{-3}$ and under uniform cost. The dashed line in each subplot represent fixed iteration stopping (e.g., the y-axis value of the line at iteration $50$ represent the cost-adjusted regret when always stopping at iteration $50$).
@@
Figure 22: Comparison of cost-adjusted simple regret across acquisition function and stopping rule combinations in the 8D Bayesian regret experiments, with cost-scaling factor $\lambda=10^{-1},10^{-2},10^{-3}$ and under linear cost. The dashed line in each subplot represent fixed iteration stopping (e.g., the y-axis value of the line at iteration $50$ represent the cost-adjusted regret when always stopping at iteration $50$).
@@
Figure 23: Comparison of cost-adjusted simple regret across acquisition function and stopping rule combinations in the 8D Bayesian regret experiments, with cost-scaling factor $\lambda=10^{-1},10^{-2},10^{-3}$ and under periodic cost. The dashed line in each subplot represent fixed iteration stopping (e.g., the y-axis value of the line at iteration $50$ represent the cost-adjusted regret when always stopping at iteration $50$).

### v2

@@ D.3 Additional Experiment Results: Empirical
Astudillo et al. (2021, Proposition 2) proposed modeling unknown cost $c(x)$ via
$\mathbb{E}[1/c(x)]^{-1}=\exp(\mu_{\ln c}(x)-(\sigma_{\ln c}(x))^{2}/2).$ (23)
An alternative is
$\mathbb{E}[c(x)]=\exp(\mu_{\ln c}(x)+(\sigma_{\ln c}(x))^{2}/2).$ (24)
The difference in sign before the variance term reflects how each formulation handles predictive uncertainty: (23) encourages more exploration than (24). For PBGI under the unknown-cost setting, it is more natural to replace $c(x)$ in (5) with $\mathbb{E}[c(x)]$ using (24), as this aligns with how costs enter the root-finding problem. For LogEIPC, both variants are possible—we refer to the (23) version as LogEIPC-inv and the (24) version as LogEIPC-exp. However, equivalence between PBGI and LogEIPC stopping rules and our theoretical guarantees hold only with (24) but not with (23). Accordingly, we use (24) for both methods in our experiments to maintain consistency and preserve this equivalence. Figure˜19 shows performance of acquisition function and stopping rule combinations under the unknown-cost setting, which are qualitatively similar to the known-cost setting.
Figure 19: Comparison of cost-adjusted simple regret across acquisition function and stopping rule combinations under the unknown-cost setting on LCBench with $\lambda=10^{-4}$. Our PBGI/LogEIPC stopping rule remains close to the hindsight optimal when paired with the LogEIPC-exp or PBGI acquisition function, sometime slightly worse then heuristics such as GSS and Convergence. It is also slightly worse than Immediate on Adult, probably due to a cost-model mismatch under the unknown-cost setting.
D.4 Additional Experiment Results: Bayesian Regret
@@
In this section, we present the complete Bayesian regret results. Figures˜20, 21 and 22 show the 1D experiments, and Figures˜23, 24 and 25 show the 8D experiments. Each figure corresponds to one cost setting (uniform, linear or periodic) and three values of the cost-scaling factor, $\lambda=10^{-1},10^{-2},10^{-3}$. In all of the experimental results, we observe that PBGI/LogEIPC acquisition function + PBGI/LogEIPC stopping achieves cost-adjusted regret that is not only competitive with the baselines, but is also competitive regarding the best in hindsight fixed iteration stopping and often competitive even comparing to hindsight optimal stopping. These results indicate that our automatic stopping rule can replace manual selection of stopping times without loss in performance. Figures˜20, 21, 22, 23, 24 and 25 also show that our PBGI/LogEIPC stopping rule outperforms other baselines when the cost-scaling factor $\lambda$ is large, indicating that it’s an especially suitable stopping criteria when evaluation is expensive.
@@
Figure 20: Comparison of cost-adjusted simple regret across acquisition function and stopping rule combinations in the 1D Bayesian regret experiments, with cost‐scaling factor $\lambda=10^{-1},10^{-2},10^{-3}$ and under uniform cost. The dashed line in each subplot represent fixed iteration stopping (e.g., the y-axis value of the line at iteration $50$ represent the cost-adjusted regret when always stopping at iteration $50$).
@@
Figure 21: Comparison of cost-adjusted simple regret across acquisition function and stopping rule combinations in the 1D Bayesian regret experiments, with cost‐scaling factor $\lambda=10^{-1},10^{-2},10^{-3}$ and under linear cost. The dashed line in each subplot represent fixed iteration stopping (e.g., the y-axis value of the line at iteration $50$ represent the cost-adjusted regret when always stopping at iteration $50$).
@@
Figure 22: Comparison of cost-adjusted simple regret across acquisition function and stopping rule combinations in the 1D Bayesian regret experiments, with cost‐scaling factor $\lambda=10^{-1},10^{-2},10^{-3}$ and under periodic cost. The dashed line in each subplot represent fixed iteration stopping (e.g., the y-axis value of the line at iteration $50$ represent the cost-adjusted regret when always stopping at iteration $50$).
@@
Figure 23: Comparison of cost-adjusted simple regret across acquisition function and stopping rule combinations in the 8D Bayesian regret experiments, with cost‐scaling factor $\lambda=10^{-1},10^{-2},10^{-3}$ and under uniform cost. The dashed line in each subplot represent fixed iteration stopping (e.g., the y-axis value of the line at iteration $50$ represent the cost-adjusted regret when always stopping at iteration $50$).
@@
Figure 24: Comparison of cost-adjusted simple regret across acquisition function and stopping rule combinations in the 8D Bayesian regret experiments, with cost‐scaling factor $\lambda=10^{-1},10^{-2},10^{-3}$ and under linear cost. The dashed line in each subplot represent fixed iteration stopping (e.g., the y-axis value of the line at iteration $50$ represent the cost-adjusted regret when always stopping at iteration $50$).
@@
Figure 25: Comparison of cost-adjusted simple regret across acquisition function and stopping rule combinations in the 8D Bayesian regret experiments, with cost‐scaling factor $\lambda=10^{-1},10^{-2},10^{-3}$ and under periodic cost. The dashed line in each subplot represent fixed iteration stopping (e.g., the y-axis value of the line at iteration $50$ represent the cost-adjusted regret when always stopping at iteration $50$).